what is the centroid of x^2=y, x+y=6, x=0. with solution


1
Expert's answer
2022-02-14T17:29:59-0500

Solution:


"x^2=y, x+y=6
\\ \Rightarrow f(x)=x^2, g(x)=6-x"

"a=-3,b=2"

Now, "A=\int_a^b[f(x)-g(x)]dx"

"A=\int_{-3}^2[x^2-6+x]dx
\\=[\dfrac {x^3}3-6x+\dfrac{x^2}2]_{-3}^2
\\=[\dfrac {8}3-12+2]-[-9+18+\dfrac{9}2]
\\=-\dfrac{125}6
\\=\dfrac{125}6" [neglecting negative sign as it is area]

Now, "\bar x=\dfrac 1A\int_a^bx[f(x)-g(x)]dx"

"\bar x=\dfrac {6}{125}\int_{-3}^2x[x^2-6+x]dx
\\=\dfrac {6}{125}\int_{-3}^2[x^3-6x+x^2]dx
\\=\dfrac {6}{125}[\dfrac{x^4}4-3x^2+\dfrac{x^3}3]_{-3}^2
\\=\dfrac{6}{125}\cdot \dfrac{125}{12}
\\=\dfrac{1}{2}"

"\bar y=\dfrac 1A\int_a^b(\dfrac{f(x)+g(x)}2)[f(x)-g(x)]dx
\\=\dfrac {6}{125}\int_{-3}^2(\dfrac{x^2+6-x}2)[x^2-6+x]dx
\\=\dfrac {3}{125}\int_{-3}^2(x^4-x^2+12x-36)dx
\\=\dfrac {3}{125}[\dfrac{x^5}{5}-\dfrac{x^3}{3}+6x^2-36x]_{-3}^2
\\=\dfrac{6}{125}\left(-\dfrac{500}{3}\right)
\\=-8"

Thus, centroid is "(\dfrac12,-8)".